OpenStudy (anonymous):

Use the half angle formulas to simplify the expression. -sqrt ((1+cos 8x)/(1-cos 8x))

OpenStudy (campbell_st):

well can't you use \[\cos^2(4x) = \frac{1 + \cos(8x)}{2}~~~or~~~2\cos^2(4x) = 1 + \cos(8x)\] and \[\sin^2(4x) = \frac{1 - \cos(8x)}{2}~~~~or~~~~2\sin^2(4x) = 1 - \cos(8x)\] sustitute them and you have \[- \sqrt{\frac{2\cos^2(4x)}{2\sin^2(4x)}}\]

OpenStudy (campbell_st):

and then you should be able to work from there

OpenStudy (anonymous):

how do you know theyre equal to that?

OpenStudy (anonymous):

Not really Im so confused.

OpenStudy (anonymous):

Yes, its the half angle one im confused with

OpenStudy (anonymous):

Why would I be using the double angle formulas?

OpenStudy (anonymous):

I have them I just dont know what to use

OpenStudy (anonymous):

So I need to use double angle formulas and not the half angle ones?

OpenStudy (anonymous):

ok where do i plug in the expression into the formula

OpenStudy (campbell_st):

i"ll eliminate all posts so that the information is in sequence... and see if it makes sense. and \[-\sqrt{\frac{2\cos^2(4x)}{2\sin^2(4x)}} = -\sqrt{\cot^2(4x)}\] just simplify from here